Status Report

NASA ISS On-Orbit Status 9 June 2009

By SpaceRef Editor
June 10, 2009
Filed under , , ,
NASA ISS On-Orbit Status 9 June 2009

All ISS systems continue to function nominally, except those noted previously or below. Another short day for the Exp-20 crew due to the early Orlan EVA-23 egress into the SM PkhO tomorrow morning (2:45am EDT).

EVA-23 sleep cycle shift:

  • Wake today – 2:00am; sleep – 10:00am EDT;
  • Wake tonight – 7:00pm; sleep – 8:20am (6/10);
  • Next wake – 2:00am (6/11).

FE-1 Barrat & FE-2 Wakata continued their new recording rounds for the experiment SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight) logging data from their Actiwatch to the HRF-1 (Human Research Facility 1) laptop as part of a week-long session. It is the third for Mike, the fourth for Koichi. [To monitor the crewmember’s sleep/wake patterns and light exposure, Mike & Koichi wear a special Actiwatch device which measures the light levels encountered by them as well as their patterns of sleep and activity throughout the Expedition and use the payload software for data logging and filling in questionnaire entries in the experiment’s laptop session file on the HRF-1 laptop. The log entries are done within 15 minutes of final awakening for seven consecutive days.]

CDR Padalka & FE-1 Barratt reviewed the timeline for tomorrow’s EVA-23, a 37-min excursion into the evacuated SM PkhO (Service Module Transfer Compartment) for installing a docking cone in place of the current flat hatch cover on the PkhO docking port. They also reviewed the uplinked depress/repress procedure for the PkhO, both reviews supported by ground specialist tagup.

Preparations by Gennady & Mike then included –

  • Preparation of sealing tools & hatch restraints,
  • Readying two sealing handles,
  • Establishing procedure for securing the covers in the PkhO with rubber retention devices,
  • Reviewing what to do in case of a hatch leak,
  • Preparing three “Pille-MKS” radiation dosimeters, recording their dosages and equipping each of the two Orlan-MK suits (in pocket on left calf) with a sensor unit (A0309 & A0310), with the third sensor remaining on the Pille Reader tray to take SM background readings, and
  • Preparing the usual medkit for the EVA-23.

FE-2 Wakata meanwhile continued preparations for the STS-127/2JA spacewalks, specifically –

  • Inspecting base mount & clutch screws of the BRT (Body Restraint Tether),
  • Stowing U.S EVA tools no longer needed after the recent Orlan EVA-22, and
  • Completing Part 1 of configuring 2J/A EVA tools (this work being scheduled in parts on three days between now and 2J/A).

FE-3 Romanenko had 2 hrs set aside for conducting an audit/inspection of all the hardware connected to RS (Russian Segment) power outlets, focusing on Russian & US plug-ins in the SM, power outlets in the DC1 (Docking Compartment), FGB and additional plug-ins powered from the Russian PPS-31 & PPS-26 panels.

In the Kibo JPM (JEM Pressurized Module), FE-4 Thirsk supported JAXA ground-controlled payload work (FACET) by activating the PLT (Payload Laptop Terminal).

Working in the U.S. WHC (Waste & Hygiene Compartment), FE-5 DeWinne removed the pretreat tank (#805028) and replaced it with a new unit (#805030).

Due to today’s maneuver of the SSRMS (Space Station Remote Manipulator System) into the viewing field of the Lab science window (starting at ~9:00am), Bob Thirsk closed the protective shutters of the window, deactivating the AgCam (Agricultural Camera) system by necessity.

After setting up the video equipment for recording the subsequent activity, Frank DeWinne & Bob Thirsk (Subject) completed the PFE protocol, a monthly 1.5-hr. procedure which checks up on blood pressure and electrocardiogram (ECG) during programmed exercise on the CEVIS (Cycle Ergometer with Vibration Isolation) in the Lab. Frank assisted the FE-4 as CMO (Crew Medical Officer). Readings were taken with the BP/ECG (blood pressure/electrocardiograph) and the HRM (heart rate monitor) watch with its radio transmitter. Afterwards, the video gear was stowed again. [BP/ECG provides automated noninvasive systolic and diastolic blood pressure measurements while also monitoring and displaying accurate heart rates on a continual basis at rest and during exercise. During previous PFE sessions, issues have been reported with the BP cuff ops. The BP/ECG requires good ECG data to operate and record an accurate BP reading. Correct placement of the microphone and securely donning the BP cuff will aid in this. The video coverage is required for biomechanical evaluation of the on-orbit crewmember, and evaluation of the on-orbit setup of equipment during data collection.]

With the first in-flight session of the ICV (Integrated Cardiovascular) assessment plus another session of the CCISS (Cardiovascular Control on Return from the ISS) experiment approaching for Thirsk & DeWinne, the FE-4 & FE-5 reviewed an uplinked Overview of the protocol, then took up reference material for their respective roles (FE-4 starting out as the Subject, FE-5 as the Operator), and tagged up with the ground team to discuss the activities ahead. The first ICV session will include Echo or Ambulatory Monitoring. [Goal of the ICV experiment is to quantify the extent, time course, and clinical significance of cardiac atrophy and identify its mechanisms. The primary objective of the accompanying CCISS experiment is to maximize the information about changes in cardiovascular and cerebrovascular function that might compromise the ability of astronauts to meet the challenge of return to an upright posture on Earth.]

Frank also performed the periodic WRS (Water Recovery System) sample analysis in the TOCA (Total Organic Carbon Analyzer), after first priming (filling) the TOCA water sample hose. After the approximately 2 hr TOCA analysis, results were transferred to SSC-7 (Station Support Computer 7) via USB drive for downlink, and the data were also logged for calldown. [The current procedure is a work-around for TOCA’s failed catalyst.]

Roman performed another inspection of the 4GB4 Hydraulic Unit of the KOB-2 (Loop 2) of the Russian SOTR Thermal Control System, checking for presence of coolant. [On 5/19, the CDR had replaced a pump unit of the 4SPN1 replaceable pump panel at this location.]

Afterwards, the FE-3 performed routine inspection & maintenance (filter cleaning) on the Russian SRVK condensate water processor of the SVO water supply system.

Later, Romanenko completed the routine daily servicing of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM. [Regular daily SOZh maintenance consists, among else, of checking the ASU toilet facilities, replacement of the KTO & KBO solid waste containers and replacement of EDV-SV waste water and EDV-U urine containers.]

The FE-3 also conducted the daily IMS (Inventory Management System) maintenance, updating/editing its standard “delta file” including stowage locations, for the regular weekly automated export/import to its three databases on the ground (Houston, Moscow, Baikonur).

CDR Padalka performed the frequent status check on the Russian BIO-5 Rasteniya-2 ("Plants-2") experiment, verifying proper operation of the BU Control Unit and MIS-LADA Module fans (testing their air flow by hand). Gennady also took photos of the plants in the LADA greenhouse using the Nikon D2

SpaceRef staff editor.